# Gravitational Wave Probe of Gravitational Dark Matter from Preheating

### v2

@@ Abstract
We forecast high-frequency gravitational wave (GW) from preheating hosting gravitational dark matter (GDM) as the indirect probe of such GDM. We use proper lattice simulations to handle resonance, and to solve GW equation of motion with the resonance induced scalar field excitations as source term. Our numerical results show that Higgs scalar excitations in Higgs preheating model give rise to magnitudes of GW energy density spectra of order $10^{-10}$ at frequencies $10-10^{3}$ MHz depending on the GDM mass, whereas inflaton fluctuation excitations in inflaton self-resonant preheating model yield magnitudes of GW energy density spectrum up to $10^{-9}~{}(10^{-11})$ at frequencies near $30~{}(2)$ MHz for the index $n=4~{}(6)$ with respect to the GDM mass of $1.04~{}(2.66)\times 10^{14}$ GeV.
@@ 2.1 Lattice treatment on parameter resonance
we solve eq.(2.1) in terms of publicly available lattice code CosmoLattice [42, 43], by adopting the lattice size $N=128^{3}$ and the minimal infrared cut-off $k_{IR}=1$ in the following discussions of this section.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 2.2 Gravitational wave production
In terms of the GW spectra at the end time of stimulation, with $a(t_{*})/a_{\rm{end}}=45$ as shown in fig.2, we show the present-day GW spectra $\Omega_{\rm{GW},0}(f)$ with respect to various values of $\kappa$ in the left panel of fig.3, by evaluating $\Omega_{\rm{GW}}(k,t_{*})$ from $t_{*}$ to the present time $t_{0}$ as shown in the appendix.A. In this figure the dependence of $\Omega_{\rm{GW},0}(f)$ on the model parameter $\kappa$ has been replaced by that of $\Omega_{\rm{GW},0}(f)$ on the GDM mass $m_{\chi}$, using the one-to-one correspondence between $\kappa$ and $m_{\chi}$ [18]. As $\kappa$ (alternatively $m_{\chi}$) increases, the magnitude of $\Omega_{\rm{GW},0}(f)$ is larger in the large $f$ regime, but it is smaller in the small $f$ regime, as previously seen in the dependence of $P_{\tilde{h}_{\lambda}}(k)$ on $\kappa$ in fig.1. Compared to the parameter resonance induced GWs studied by [31, 33], the Higgs resonance induced GWs spectra in fig.3 have relatively smaller peak frequencies and larger peak energy densities, as the values of $q$ defined by [33] in the Higgs preheating have been constrained into a smaller range of $q\sim 337-6337$.
@@ 3 The minimal preheating
with the index $n$ being an even number in order to make sure the positivity of the inflation potential. As explained in the beginning of Sec.2, the model parameter $\lambda_{\phi}$ has been fixed by the Planck data, implying that $n$ is the only free parameter in eq.(13).
@@ 3.1 Lattice treatment on self-resonance
Using the initial conditions displayed in Table 1 of [19] and adopting the lattice size $N=256^{3}~{}(128^{3})$ and the minimal infrared cut-off $k_{IR}=0.2~{}(5\times 10^{-4})$ for $n=4~{}(6)$ [^2] One has to choose different values of $k_{IR}$ in order to capture the different resonant bands with respect to different values of $n$. to solve eq.(3.1), we show the evolution of $P_{\tilde{\delta\phi}}(k)$ with respect to time in fig.4, with $k$ in units of $\omega_{*}$. The explicit values of $\omega_{*}$ can be found in [19]. Similar to the parameter resonance induced Higgs power spectrum $P_{\tilde{h}}(k)$ in Sec.2, the self-resonance induced power spectrum $P_{\tilde{\delta\phi}}(k)$ in this figure rapidly grows until the backreaction becomes strong. Afterward, it becomes stabilized as shown by the red curves with respect to $a(t_{*})/a_{\rm{end}}\approx 1.2\times 10^{3}~{}(1.1\times 10^{5})$ for $n=4~{}(6)$ respectively.
@@ 3.2 Gravitational wave production
Using the results of fig.5 we show the present-day GW spectra in fig.6, by evaluating $\Omega_{\rm{GW}}(k)$ from the end time of preheating $t_{*}$ to the present time $t_{0}$ as shown in Appendix.A. In this figure several local peak frequencies previously seen in fig.5 appear near $\sim 40~{}(1)$ MHz for $n=4~{}(6)$. Moreover, the peak value of $\Omega_{\rm{GW},0}$ for $n=4$ is about two orders of magnitude larger than that for $n=6$, due to the relatively larger inflaton fluctuation excitations as seen in fig.4. Our results in fig.6 differ from those in the literature. First, GWs reported by refs.[32, 34, 35] are based upon an inflation potential being different from our choice in eq.(14). Second, GWs in ref.[37] arise from a broad resonance with the mass scale in eq.(14) far less than $M_{P}$ instead of a narrow resonance considered here. Lastly, while focusing on the narrow resonance, ref.[41] has only considered GW with respect to the case of $n=4$.
@@ 4 Conclusion
There are several aspects left for future study. First, our idea of using GWs as the indirect probe of GDM from preheating can be also applied to GDM from reheating, despite more diverse GW sources expected in the later situation. Second, the dependences of GW spectra on the lattice size $N$ and cut-off $k_{IR}$ can be further suppressed by more powerful compute sources. Lastly, in order to eliminate the uncertainty in deriving the present-day GW spectra a more complete treatment on the equation of state between the end of preheating and of reheating than in the Appendix is needed.
Acknowledgments
We thank the referee for the valuable comments and suggestions.
@@ Appendix A Evaluating gravitational wave spectrum
where $k_{*}$ is replaced by $f$ via eq.(A) and in the second line $\Omega_{\rm{rad},0}\approx 5\times 10^{-5}$ has been used. The main uncertainty in eq.(A) arises from $\epsilon_{*}$ which is sensitive to the value of $\bar{\omega}$.

### v3

@@ Abstract
We forecast high-frequency gravitational wave (GW) from preheating hosting gravitational dark matter (GDM) as the indirect probe of such GDM. We use proper lattice simulations to handle resonance, and to solve GW equation of motion with the resonance induced scalar field excitations as source term. Our numerical results show that Higgs scalar excitations in Higgs preheating model give rise to magnitudes of GW energy density spectra of order $10^{-10}$ at frequencies $10-10^{3}$ MHz depending on the GDM mass of $(6-9)\times 10^{13}$ GeV, whereas inflaton fluctuation excitations in inflaton self-resonant preheating model yield magnitudes of GW energy density spectrum up to $10^{-9}~{}(10^{-11})$ at frequencies near $30~{}(2)$ MHz for the index $n=4~{}(6)$ with respect to the GDM mass of $1.04~{}(2.66)\times 10^{14}$ GeV.
@@ 2 The Higgs preheating
Within this resonant parameter range, the GDM mass $m_{\chi}\sim(4-6)m_{\phi}\sim(6-9)\times 10^{13}$ GeV due to the one-to-one correspondence [18] between $\kappa$ and $m_{\chi}$ imposed by the observed DM relic density.
@@ 2.1 Lattice treatment on parameter resonance
we solve eq.(2.1) in terms of publicly available lattice code CosmoLattice [42, 43], by adopting the lattice size $N=128^{3}$ and the minimal infrared cut-off $k_{IR}=1$ in the following discussions of this section. In this code the Hubble parameter is solved self-consistently through Friedmann equation.
@@ 2.2 Gravitational wave production
In terms of the GW spectra at the end time of stimulation, with $a(t_{*})/a_{\rm{end}}=45$ as shown in fig.2, we show the present-day GW spectra $\Omega_{\rm{GW},0}(f)$ with respect to various values of $\kappa$ in the left panel of fig.3, by evaluating $\Omega_{\rm{GW}}(k,t_{*})$ from $t_{*}$ to the present time $t_{0}$ as shown in the appendix.A. In this figure the dependence of $\Omega_{\rm{GW},0}(f)$ on the model parameter $\kappa$ has been replaced by that of $\Omega_{\rm{GW},0}(f)$ on the GDM mass $m_{\chi}$, using the one-to-one correspondence between $\kappa$ and $m_{\chi}$ [18]. As $\kappa$ (alternatively $m_{\chi}$) increases, the magnitude of $\Omega_{\rm{GW},0}(f)$ is larger in the large $f$ regime, but it is smaller in the small $f$ regime, as previously seen in the dependence of $P_{\tilde{h}_{\lambda}}(k)$ on $\kappa$ in fig.1. Compared to the parameter resonance induced GWs studied by [31, 33], the Higgs resonance induced GWs spectra in fig.3 have relatively smaller peak frequencies and larger peak energy densities.
@@ 3 The minimal preheating
with the index $n$ being an even number in order to make sure the positivity of the inflation potential. As explained in the beginning of Sec.2, the model parameter $\lambda_{\phi}$ has been fixed by the Planck data, implying that $n$ is the only free parameter in eq.(13). For the specific situation with $n=4~{}(6)$, the GDM mass $m_{\chi}\approx 1.04~{}(2.66)\times 10^{14}$ GeV due to the one-to-one correspondence [19] between $m_{\chi}$ and $n$ imposed by the observed DM relic density.
@@ 3.1 Lattice treatment on self-resonance
Using the initial conditions displayed in Table 1 of [19] and adopting the lattice size $N=256^{3}~{}(128^{3})$ and the minimal infrared cut-off $k_{IR}=0.2~{}(5\times 10^{-4})$ for $n=4~{}(6)$ [^2] One has to choose different values of $k_{IR}$ in order to capture the different resonant bands with respect to different values of $n$. to solve eq.(3.1), we show the evolution of $P_{\tilde{\delta\phi}}(k)$ with respect to time in fig.4, with $k$ in units of $\omega_{*}$. The explicit values of $\omega_{*}$ can be found in [19]. Similar to the parameter resonance induced Higgs power spectrum $P_{\tilde{h}}(k)$ in Sec.2, the self-resonance induced power spectrum $P_{\tilde{\delta\phi}}(k)$ in this figure rapidly grows until the backreaction becomes strong. Afterward, it becomes stabilized as shown by the red curves with respect to $a(t_{*})/a_{\rm{end}}\approx 1.2\times 10^{3}~{}(1.1\times 10^{5})$ for $n=4~{}(6)$ respectively. Compared to $n=4$ in the left plot, the resonant peaks for $n=6$ in the right plot are much more suppressed. The reason is that the oscillating frequency $\omega_{k}=(k^{2}+a^{2}\partial^{2}V/\partial\phi^{2})^{1/2}\sim a^{0}$ and $\sim a^{-1/2}$ for $n=4$ and 6 respectively, implying a relatively postponed growth of the inflaton fluctuations in the later case.
@@ 3.2 Gravitational wave production
Using the results of fig.5 we show the present-day GW spectra in fig.6, by evaluating $\Omega_{\rm{GW}}(k)$ from the end time of preheating $t_{*}$ to the present time $t_{0}$ as shown in Appendix.A. In this figure several local peak frequencies previously seen in fig.5 appear near $\sim 40~{}(1)$ MHz for $n=4~{}(6)$. Moreover, the peak value of $\Omega_{\rm{GW},0}$ for $n=4$ is about two orders of magnitude larger than that for $n=6$, due to the relatively larger inflaton fluctuation excitations as seen in fig.4. Compared to our results, (i) GWs in refs.[32, 34, 35] are based upon an inflation potential being different from our choice in eq.(14); (ii) GWs in ref.[37] arise from a broader resonance with the mass scale in eq.(14) less than $M_{P}$ considered here; and (iii) Ref.[41] has only considered the GW for $n=4$, where the peak values of $k/m_{\rm{end}}\approx\{1.0,1.5,2.0,\cdots\}$ with $m_{\rm{end}}=\sqrt{3}\omega_{*}$ correspond to the peak values of $k/\omega_{*}\approx\{1.8,2.7,3.5,\cdots\}$ in the left plot of fig.5 respectively.
@@ 4 Conclusion
There are several aspects left for future study. First, our idea of using GWs as the indirect probe of GDM from preheating can be also applied to GDM from reheating, although more diverse GW sources are expected in the later situation. Second, the dependences of GW spectra on the lattice size $N$ and cut-off $k_{IR}$ can be further suppressed by more powerful compute sources. Lastly, in order to eliminate the uncertainty in deriving the present-day GW spectra a more complete treatment on the equation of state between the end of preheating and of reheating than in the Appendix is needed.
@@ Appendix A Evaluating gravitational wave spectrum
where $k_{*}$ is replaced by $f$ via eq.(A) and in the second line $\Omega_{\rm{rad},0}\approx 5\times 10^{-5}$ has been used. The main uncertainty in eq.(A) arises from $\epsilon_{*}$ which is sensitive to the value of $\bar{\omega}$. Take $a_{*}/a_{\rm{rad}}\approx 10^{-5}$ for example, $\bar{\omega}=0.3~{}(0.32)$ gives $\epsilon_{*}=0.32~{}(0.63)$.